### v1

# Modeling and Optimizing User Preferences in AI Copilots: A Comprehensive Survey and Taxonomy

Abstract
AI copilots, context-aware, AI-powered systems designed to assist users in tasks such as software development and content creation, are becoming integral to modern workflows. As these systems grow in capability and adoption, personalization has emerged as a cornerstone for ensuring usability, trust, and productivity. Central to this personalization is preference optimization: the ability of AI copilots to detect, interpret, and align with individual user preferences. While personalization techniques are well-established in domains like recommender systems and dialogue agents, their adaptation to interactive, real-time systems like AI copilots remains fragmented and underexplored. This survey addresses this gap by synthesizing research on how user preferences are captured, modeled, and refined within the design of AI copilots. We introduce a unified definition of AI copilots and propose a phase-based taxonomy of preference optimization strategies, structured around pre-interaction, mid-interaction, and post-interaction stages. We analyze techniques for acquiring preference signals, modeling user intent, and integrating feedback loops, highlighting both established approaches and recent innovations. By bridging insights from AI personalization, human-AI collaboration, and large language model adaptation, this survey provides a structured foundation for designing adaptive, preference-aware AI copilots. It offers a holistic view of the available preference resources, how they can be leveraged, and which technical approaches are most suited to each stage of system design.
Keywords Preference Detection, Personalization, Personalized Response Generation, Cognitive Assistants, AI Copilots, Human-Centric Evaluation, Human–Computer Interaction (HCI)
1 Introduction
AI-powered, context-aware systems—commonly referred to as AI copilots—are rapidly emerging as collaborative assistants for developers, analysts, and knowledge workers engaged in complex, high-context tasks. These systems are integrated into user workflows and designed to understand evolving goals, adapt to individual preferences, and provide real-time, personalized support. However, achieving meaningful personalization in such systems presents unique challenges. User intent is often ambiguous or changes over time, tasks may span multiple domains, and responses must be dynamically adapted to nuanced contexts. These systems must be capable of aligning with dynamic user preferences, as personalization is a key success factor for digital assistants, enabling them to adapt their responses and behavior to individual users’ routines and context [1].
At the core of this personalization lies preference optimization—the ability of a system to detect, interpret, and align with user preferences [2]. While advances in preference modeling have gained traction in machine learning and recommender systems, their integration into real-time, interactive systems like AI copilots remains limited and fragmented. Designing copilots that can continuously adapt to evolving preferences in high-context environments is still a major open challenge.
This survey is driven by the need to systematically examine how preference optimization is realized in AI copilots. Our aim is to consolidate disparate research across assistant technologies, personalization, and alignment learning, and to provide a structured account of how preferences are acquired, modeled, and refined in these systems. In doing so, we offer a comprehensive view of the emerging design space for adaptive, user-centered AI copilots.
Research related to this survey broadly falls into two categories: (i) AI systems and intelligent assistants, and (ii) preference optimization techniques. While each has been explored independently in the literature, their integration remains underexamined. The evolution of virtual assistants such as Google Assistant, Siri, Cortana, and Alexa has been well documented, with studies focusing on their core functionalities, user experience, and limitations in areas such as voice recognition, contextual awareness, and multi-turn dialogue management [3]. In enterprise contexts, digital assistants have been examined through the lens of business integration, workflow support, and human-computer collaboration, highlighting both their organizational potential and associated risks [1]. Concerns around privacy and security in these systems have also been widely addressed, particularly in relation to data handling and unauthorized access [4]. Although the design and evaluation of AI copilots in retail has been recently analyzed through Microsoft’s case studies, the work is limited in scope and does not offer a comprehensive perspective on the broader cognitive assistant landscape [5]. On the other hand, preference optimization—especially in the context of aligning AI behavior with user intent—has seen significant theoretical and practical development. A recent comprehensive survey on Direct Preference Optimization (DPO) presents an in-depth analysis of its theoretical foundations, training methodologies, benchmark datasets, and diverse application scenarios [2]. However, such works primarily examine preference modeling in isolation and do not address its integration within interactive systems like AI copilots or digital assistants.
While existing studies contribute significantly to our understanding of intelligent assistant technologies and preference optimization methods, their **integration in the context of AI copilots remains limited and underexplored**. As a result, there is no consolidated view of how preference signals can be effectively incorporated into the behavior of real-time, context-aware systems. This survey aims to bridge these two streams of research by examining preference optimization techniques within the design of AI copilots—systems that operate collaboratively with users in high-context, evolving environments. Specifically, we address how preferences can be detected, modeled, and refined to enhance personalization and system responsiveness. The main contributions of this work are summarized as follows:
  • A unified and literature-grounded definition of AI copilots is proposed, synthesizing diverse descriptions and terminology from recent studies across domains.
  • A comprehensive analytical review of preference optimization in AI copilot design is provided, covering possible sources of user preferences, techniques for preference-aware response generation, and methods for feedback-driven adaptation.
We organized the content of our survey as follows: Section 2 provides foundational background by examining the historical development and evolving terminology associated with AI copilots. Section 3 introduces a unified, literature-informed definition of AI copilots, integrating insights from diverse fields. Section 4 presents a conceptual AI copilot architecture focused on preference optimization, offering a structured analysis of core components, including preference sources, detection techniques, personalization strategies, and feedback-driven refinement. Finally, Section 5 concludes with a summary of key insights and implications for future research.
2 Priliminary and backgrounds
To establish a clear foundation for defining the new generation of AI-powered systems—referred to as AI copilots—we analyze key concepts and terminology from existing literature. This section provides a precise understanding of how AI copilots differ from prior intelligent systems in function, autonomy, interaction, and domain focus. Table 1 summarizes these distinctions to support a consistent and rigorous discussion throughout the paper.
2.1 AI Copilot Related Terms
2.1.1 Cognitive Assistants
One of the concepts related to the scope of our survey in the literature is AI Assistant, which shouldn’t be used interchangeably with AI Copilot. As a general definition of AI Assistant, Danaher [6] defines it as a program that tries to optimize a required target output based on the possible range of options. This notion highlights the nature of AI assistants as an optimization tool rather than as collaborative enhancers of human abilities. As a more recent idea, AI co-pilots aim to work alongside people to improve their decision-making and capacity in more complex scenarios. The AI Co-Pilot concept, as highlighted in Sellen et al. (2023)[7], is defined as a collaborative partner aiding humans, rather than just a tool. It emphasizes assisting humans with insightful support in professional and industrial settings, enhancing the human-AI partnership with its context-aware capabilities. Therefore, while AI Assistants serve the broader consumer base, AI Co-Pilots are specifically designed with developers in mind as their target user group.
Starting with Cognitive Computing, this broad term encompasses hardware or software systems designed to mimic human cognitive abilities, as explored by Kelly (2015)[8]. In such a context, Metacognition described by Schmorrow (2005)[9] as a nuance expansion, extends the paradigm by fostering a symbolic, closed-loop integration of humans and machines. In such a continuous feedback loop, the communication channel for information sharing and decision-making is done by cognitive functions. Delving into a practical application, the Joint Cognitive System, highlighted by Hollnagel (2005)[10], presents an interactive environment where goal-oriented human and machine elements collaboratively share and integrate their cognitive abilities within specific contexts. Rooted in a human-centric approach, this strategy accentuates the synergy of human creativity with AI’s analytical capabilities, ensuring adaptability across scenarios. Cognitive assistants stand at the forefront of Human-AI Collaboration, merging human insight with AI’s vast processing capacity to improve problem-solving and decision-making [11]. Through studies in hybrid intelligence systems and empathic AI-supported conversations, these technologies demonstrate their capacity to enhance teamwork, innovation, and human empathy [12, 13]. This synergy underscores the significant benefits of blending AI with human insights, positioning cognitive assistants as key facilitators of effective human-AI partnerships.
2.1.2 Digital Assistant
The term Digital Assistant (DA) is frequently interchanged with Virtual Assistant (VA), yet it represents a broader concept. In a study about its anatomy, Wells et al. [14] referred to it as a versatile tool and amorphous compound technology. Mobility, rapid data analysis capability and supporting multiple interfaces are among its key features making it ideal for task delegation and advice provision applications. The study by Knote et al. [15] highlights a key feature of Digital Assistants (DAs): their ability to merge the simplicity of intuitive natural language dialogue with personalized, context-specific assistance. This study focuses on categorizing the different forms of the Smart Personal Assistant (SPA), which is essentially a personalized Digital Assistant. Thus, the classification of SPAs into text-based systems, often termed Chatbots, and voice-based systems, as conducted by Cowan et al. [16], can be applied to DAs. By considering Chatbots as a form of Digital Assistant, all the categories of employed technologies in their construction can be categorized under DA. This allows DA’s boundaries to be expanded to cover a wider range of technologies, from simple rule-based Chatbots with limited capabilities to intelligent voice-based VAs with speech recognition capability [17].
Table 1: Comparison of AI Copilot-Related Terms and Technologies
System Type Primary Function Level of Autonomy Interaction Modality Domain Specialization
Cognitive Assistant Advanced decision support through reasoning and learning Proactive / Semi-autonomous Multimodal (text, voice, context-aware UI) Often domain-specific (e.g., healthcare, law)
Digital Assistant Task management, scheduling, basic information retrieval Semi-autonomous Voice + UI General-purpose
Virtual Assistant Conversational help and simple task execution Reactive to semi-autonomous Voice, text General-purpose
Chatbot Text-based conversational responses Reactive Text Often domain-specific (e.g., customer support, retail)
Expert System Rule-based decision-making using encoded expert knowledge Low autonomy (deterministic) Menu-driven or text-based Highly domain-specific
AI Copilot Context-aware collaboration on complex or creative tasks Proactive Multimodal (text, code, UI, APIs) Task- or profession-specific (e.g., software development, writing)
2.1.3 Virtual Assistant
As a good transitional definition from DA to VA, Apple Siri and Amazon Alexa as two popular intelligent assistants referred to as voice-based DAs [1, 18]. Their ability to perform hands-free operations, context understanding and handle complex conversations are among the key features mentioned by Wellsandt et.al study [14]. Besides their speech recognition ability to understand voice commands and communicate over the same channel [4], voice-activation ability made them suitable for integrating into daily use devices [19]. Although they are mainly referred to as VAs, Smart Personal Assistant (SPA) is another term used for Apple Siri and Amazon Alexa as two device-integrated assistants focusing on their personalized assisting[15]. This positions them within the broader spectrum of Digital Assistants, primarily known for their voice-based interfaces, but also encompassing text-based interaction as an additional feature, distinguishing another subset of such assistants [20].
2.1.4 Chatbot
Chatbots are specialized programs tailored for customer service, helping users find desired information [21]. This focused definition underscores their application and emphasizes their role, setting them apart from other digital assistants despite shared technologies and capabilities. A defining feature of chatbots, as highlighted in literature [22], is simulating near-human conversations using natural language or text, to create an impression of engaging in dialogue with a human [23]. Emerging from the field of Human-Computer Interaction (HCI), which centres on enabling users to communicate with computers, Dialogue Systems is a term often used to describe Chatbots [21, 24]. Conversational Systems, Conversational Agents, and Chatter-bot are among other terms used interchangeably to refer to the same term [22]. While there are different criteria for Chatbot classification in the literature, they are primarily classified by their applications, falling into either Task-oriented or Non-task-oriented (open-ended or simple chit-chat) categories [21, 23].
Task-oriented chatbots generally consist of domain-specific hand-crafted rules with limited and focused conversation context analysis and are typically utilized in devices for convenience. Non-task-oriented or informative Chatbots, often handle random dialogue contexts without aiming for a specific goal and may require online search capabilities to operate effectively [24, 23]. Based on the Knowledge Domain, chatbots can be categorized into closed-domain or open-domain, where the open-domain modeling, a challenging field, is predominantly addressed using Neural Network (NN) approaches, including both Retrieval-based and Generative (Generation-based) Chatbots [22]. Retrieval-based chatbots support immediate response generation in e-business by ranking candidate responses based on similarity measurements between the query and knowledge base of QA pairs. Data-driven Generative Chatbots utilize Natural Language Processing (NLP) to extract content from queries and learn from entire dialogue contexts, rather than solely focusing on the current turn. These chatbots apply both supervised methods, which rely on annotated data and manual feature extraction, and unsupervised methods, which automatically learn features from unlabeled data. Responses are generated word by word using neural network techniques such as RNN, LSTM, and Deep Seq2Seq, all of which operate based on a learned vocabulary [22]. In this regard, RAG-based approaches as a more recent approach, combine the previous techniques for creating a better open-domain chatbot and even injecting more domain-specific knowledge by additional training [25].
2.1.5 Expert System
As one of the old terms in cognitive computing, an Expert System (ES) is defined as a computer program that emulates the behavior of a human expert [26]. They are a branch of AI (that is a high-level term) that is used for transferring task-specific knowledge from human to computer in the way that it continuously changes and obtains new knowledge[27]. Used in problem-oriented domains including social science, cognitive science and human behaviour analysis, ES has two main parts that enable it to produce expert-like advice. As the first main component, the knowledge base, constructed by knowledge engineers using the suitable notation of language, stores the codified collected knowledge. The inference engine, guided by a suitable reasoning method chosen to ensure efficient use of the knowledge, then uses the knowledge to develop a conclusion [26]. For transparency reasons, the system can benefit from an Explanation Facility to elucidate the logic behind the conclusion [28]. Although Rule-Based Systems (RBS) or Knowledge-Based Systems (KBS) are used to refer to the same term [29, 28], others consider them as main streams of ES[27]. In the RBS, knowledge representation is a collection of IF-THEN rules. KBS as a human-centred design, includes knowledge engineering tools and a user interface for this reason. Case-Based Reasoning (CBR) systems in one category of ESs adapt solutions to new challenges by comparing them to past cases based on set parameters. Meanwhile, Fuzzy Expert Systems (Fuzzy ESs) in another category, embrace uncertainty in human reasoning, transcending binary logic for nuanced decision-making in complex situations where traditional approaches fall short [27].
The Belief Rule-Based Expert System (BRBES) merges the accuracy of expert systems with the adaptability of data-driven models, adeptly handling quantitative and qualitative data to address complexities in uncertain environments [30]. It manages uncertainties—like fuzziness, randomness, and ignorance—through belief degrees and employs the Evidence Reasoning (ER) algorithm for seamless data integration. This approach makes BRBES a transparent and dependable solution for decision-making in situations marked by scarce or ambiguous information.
3 AI Copilot Definition
Figure 1: Conceptual design architecture of a preference-aware AI copilot, illustrating the flow from user input and preference modeling to personalized response generation and feedback-driven optimization.
The term AI copilot has gained significant attention across various domains, yet its definition remains context-dependent. To establish a clear understanding, this section reviews how recent studies describe AI copilots across both general and domain-specific contexts, aiming to distill a unified definition grounded in the literature.
AI copilots have been described as intelligent assistants designed to support users in executing cognitive tasks [5]. Their architectures are typically composed of large language models, mechanisms for knowledge retrieval and action execution, orchestration modules to manage subtasks, structured system prompts, and responsible AI safeguards to ensure alignment with human intent. In domain-specific applications such as pathology, copilots have been tailored through fine-tuning on real-world instructions to address complex tasks like question answering [31]. These systems are often multimodal and interactive, capable of processing both visual and textual inputs, and are used to assist human-in-the-loop decision-making processes. In programming contexts, copilots have been implemented to provide real-time suggestions that support knowledge workers and improve productivity, while allowing users the autonomy to accept, reject, or modify the outputs [32]. A broader vision has been presented in which copilots are viewed as collaborative partners that operate alongside human users, each enhancing the other’s capabilities [33]. This metaphor has further been extended to emphasize fluid, conversational interaction and limited autonomy in pursuing shared goals under human supervision [34]. Taken together, these perspectives highlight AI copilots as interactive, task-oriented systems that support users, particularly knowledge workers, through real-time assistance, domain adaptation, and human-guided decision-making. These shared characteristics form the foundation for a unified understanding of AI copilots presented in this survey.
4 Preference-based Response Generation
Considering user preferences in an AI copilot as a system-level attribute that must be addressed across different stages. Fig. 1 shows the conceptual design of an AI copilot from a preference-aware perspective. It briefly illustrates the journey of user preferences—from input channels, through the development of preference representations, to their use in response generation and, finally, feedback-driven optimization. In the subsequent sections, we review related studies for each part of this preference-aware design in detail.
4.1 Preference Detection Sources
The effectiveness of AI copilots in aligning with user needs largely depends on the quality and nature of the preference signals they rely on. This section categorizes the primary sources of such signals, spanning both user-driven and model-generated inputs. Table 2 offers a comparative overview of their underlying mechanisms, applications, and associated limitations, providing a foundation for the analyses that follow.
Table 2: Taxonomy of Preference Detection Sources: Mechanisms, Use Cases, and Limitations
Source Category Core Mechanisms Use Cases Limitations
Explicit Feedback Pairwise comparisons (e.g., Chatbot Arena) [35] Choice prompts [36] Satisfaction vs. engagement signals [37] Direct preference labeling Evaluation benchmarking Task-specific tuning High annotation cost User fatigue Low scalability
Implicit Human Feedback Clickstream analysis [38] Gaze-based attention [39] User edits as coactive signals [40] Low-friction feedback Behavioral modeling Passive signal collection Privacy issues Sparse supervision
Hybrid Feedback Models RLHF-style fusion [41] Hashtag-enriched profiles [42] Behavioral–explicit alignment Long-term preference tracking Integration complexity Data inconsistency System dependence
Interactive Elicitation Active LLM probing [43] Bayesian acquisition [44] Probabilistic elicitation [45, 46] Uncertainty-aware querying Adaptive tuning Query design complexity Limited domain coverage
LLM-Generated Signals Simulated user modeling [47] Profile-conditioned tuning [48] Alignment via divergence minimization [49] Multi-turn preference tracking [50] Synthetic supervision Long-context adaptation Lack of grounding Hallucination bias Evaluation ambiguity
4.1.1 Human-Driven Feedback Channels
Preference detection in personalized response generation systems predominantly relies on human-generated signals, which vary from implicit behavioural cues to explicit user feedback. Implicit feedback is particularly valuable due to its non-intrusive nature, allowing preference inference without requiring active user involvement. Early research in this domain utilized web interaction data, applying web usage mining to infer user preferences from browsing behaviours [38]. Subsequent methods have evolved to integrate richer implicit signals, such as gaze tracking and physiological responses. For example, in a related study [39], gaze-tracking technology is used to predict user preferences for video content with significant accuracy, highlighting the utility of visual attention cues in multimedia contexts. Additionally, to reduce user burden while still capturing meaningful preferences, more nuanced forms of implicit feedback have been explored. In a recent study [40], a coactive learning approach was introduced, where minor user edits to AI-generated outputs are interpreted as weak preference signals. It was shown that incorporating these subtle modifications can incrementally refine language model outputs and enhance personalization without requiring extensive user input. Nonetheless, the passive collection of implicit signals raises significant privacy concerns. To address these, frameworks such as Heimdall [51] aggregate and anonymize user data securely, ensuring personalization benefits without compromising user privacy. However, implicit feedback inherently lacks granularity, motivating the integration of explicit preference elicitation strategies.
Explicit feedback approaches directly solicit user preferences, typically via ratings or comparative evaluations, providing clear and interpretable signals for personalization. Platforms such as Chatbot Arena [35] exemplify the utility of explicit feedback, employing pairwise comparisons of model-generated outputs to obtain cognitively efficient yet precise preference signals. This method has proven highly effective for benchmarking language model performance based on human preference. To mitigate limitations such as user fatigue inherent in traditional explicit methods, generative augmentation techniques have been introduced in recent work [36]. These techniques dynamically tailor choice-based prompts to increase user engagement and enhance the clarity and reliability of elicited preferences. The value of explicit feedback has been demonstrated through empirical evaluations [37], where extensive user studies revealed that satisfaction ratings outperform implicit engagement metrics in aligning with actual user preferences. These findings underscore the importance of explicit signals in achieving accurate personalization. In response, hybrid strategies have been developed to integrate the complementary strengths of both explicit and implicit feedback. For instance, topical user profiling integrates explicit declarations of interest with implicit behavioural signals (e.g., hashtag usage), significantly enhancing the robustness and reliability of personalized recommendations [42, 41]. Despite their clarity, explicit methods inherently demand continuous active user engagement, motivating the development of interactive and hybrid elicitation approaches.
Interactive and hybrid preference elicitation strategies aim to balance explicit clarity and implicit convenience, reducing user effort while maintaining high-quality feedback signals. One promising interactive method [43] involves self-exploring language models, which autonomously generate diverse output variations to probe user preferences. The effectiveness of such models has been demonstrated through reinforcement learning loops that iteratively refine responses based on real-time user reactions. This approach substantially reduces user effort while improving alignment accuracy. Furthermore, optimization-driven elicitation strategies utilizing Bayesian optimization have been proposed to systematically minimize user interactions. Specifically, LLM-based acquisition functions [44] have been introduced to strategically select the most informative queries, significantly decreasing interaction requirements without sacrificing the quality of preference inference. Complementary approaches have employed probabilistic reasoning [46] to dynamically adapt elicitation strategies based on uncertainty in user feedback, enabling effective handling of ambiguous interactions. In parallel, a Bayesian preference elicitation framework [45] has been developed to rigorously integrate prior knowledge with real-time user responses, enhancing the statistical validity and reliability of interactive preference modeling in conversational settings. Collectively, these interactive and hybrid elicitation methods represent a significant advancement in preference detection, optimizing user involvement and improving feedback clarity in personalized response generation systems.
4.1.2 LLM-Generated Feedback Channels
Complementing human-generated signals, recent advances increasingly leverage autonomous large language models (LLMs) as intrinsic sources of preference detection. Structured persona and profile modeling constitute a primary approach, wherein user contexts are systematically synthesized into structured intermediate representations prior to response generation. For example, a guided profile generation techniques is introduced [48], demonstrating a notable improvement in preference alignment accuracy compared to direct raw-context utilization, highlighting the efficacy of structured user profiling for personalized dialogue response generation. Optimization-driven methodologies offer another dimension of LLM-generated feedback, explicitly aligning model outputs with user preferences through targeted minimization of divergence metrics. For instance, an f-divergence minimization technique [49] is applied to finely control the alignment between user-preferred and model-generated responses, effectively balancing consistency and diversity in personalized outputs.
Beyond structured and optimization-driven methodologies, conversational preference elicitation frameworks systematically harness LLMs to guide user interactions, capturing rich preference signals. The Coached Conversational Preference Elicitation (CCPE) framework [52] demonstrated that structured yet naturalistic conversational interactions capture significantly richer user preferences compared to traditional static feedback mechanisms. Extending this approach, a multi-turn interaction framework [50] is presented wherein LLMs incrementally adapt to evolving user preferences through repeated conversational exchanges, resulting in improved alignment accuracy as shown on the ALOE benchmark. Additionally, the long-term retention of user preferences in LLMs has been evaluated using benchmarks like PrefEval [53], which revealed notable limitations in maintaining and applying preferences over extended interactions. These findings point to key challenges for future research in conversational personalization.
Finally, generative user simulation techniques employing LLMs themselves as realistic user models have emerged as crucial tools for scalable and systematic evaluation of personalized response generation. Generative user simulations [47] have been shown to effectively emulate dynamic shifts in user preferences over time, enabling controlled experimental environments for rigorous evaluation and iterative refinement of personalization models without relying on extensive real-world testing. Taken together, these methodologies employing LLM-generated signals offer substantial promise for enhancing the scalability, adaptability, and precision of preference detection, crucially supporting robust and dynamic personalized AI response generation frameworks.
4.2 Preference Detection Techniques
Following the identification of preference signal sources, this section turns to techniques for detecting and modeling user preferences across different phases of interaction. These are typically categorized into pre-, mid-, and post-interaction stages, each offering distinct opportunities for adaptation. A summary of representative techniques and their characteristics is presented in Table 3.
4.2.1 Predefined Profiles and Persona Development
Before any user interaction begins, preference detection is typically performed through predefined user profiles. These profiles are constructed by collecting and analyzing structured data that includes demographic information, behavioral patterns, and previously recorded interactions. These profiles allow AI systems to generate context-aware and personalized responses from the outset, enabling more efficient alignment with expected user behavior [54], [55].
The process of user profiling has undergone significant evolution. Initially, the focus was placed on static features, including age, gender, and interaction frequency. While these attributes provided a coarse approximation of user preferences, they lacked adaptability and generalization. The inability to account for contextual variation or multi-turn interaction led to the exploration of more advanced persona-aware techniques. In such approaches, user preferences were estimated before dialogue using aggregated observations and preference indicators [56]. To address the limitations of static profiling, neural models such as SessionBERT were introduced. This model incorporated session-level contextual cues to refine user representation over short sequences of interaction, even prior to any real-time dialogue [57]. Unlike earlier statistical models, SessionBERT adapted to usage behavior and allowed profile updates based on session data. By generating session embeddings and conditioning predictions on these dynamic vectors, SessionBERT improved personalization in early conversational stages. This resolved the challenge of rigidity seen in earlier demographic or rule-based systems and introduced a scalable method to encode latent preference variation [57], [56]. In parallel, collective behavior modeling was explored to capture user norms and group-aligned preferences. Methods based on collaborative filtering were applied to infer unknown preferences by observing similar users’ selections, thereby enriching persona development prior to interaction [58]. Compared to isolated profiling, these techniques introduced a population-based signal that improved personalization accuracy and generalization. In this stage, persona development has played a central role in preparing systems for adaptive interaction. It has been supported by a combination of statistical analysis, collaborative filtering, and multi-dimensional modeling of user characteristics [59]. These techniques have enabled the construction of detailed, structured user representations that anticipate needs even before conversation begins. The shift from rigid user profiles to enriched, data-driven personas has improved alignment between initial system output and user intent.
4.2.2 Real-Time Persona Extraction and Preference Adaptation
In contrast to predefined user profiles developed before interaction, the during the conversation phase focuses on dynamically identifying and adapting user preferences in real-time. In early approaches, systems were only capable of reactive adaptation, adjusting responses based on superficial turn-level cues. These methods could not infer persistent user traits or adapt to evolving intent across dialogue segments.
To address these shortcomings, neural collaborative filtering models were explored to implicitly capture preferences. One such method, SNCF (Social Neural Collaborative Filtering), was proposed to model user preferences from biased implicit feedback using neural embeddings [60]. In SNCF, user-item interaction histories were encoded into latent vectors, enabling the capture of nuanced behavioral patterns that could be refined during interactions. Although initially designed for recommendation tasks, the model’s ability to represent evolving user preferences informed later approaches to conversational preference adaptation. As the need for more fine-grained, dialogue-level understanding grew, real-time persona extraction methods were introduced. PAED enabled persona attributes to be extracted directly from conversational content without relying on external profiles [61]. By detecting triplet-style attributes within dialogue turns, PAED facilitated lightweight personalization but remained limited to surface-level lexical features. To enrich the contextual modeling of persona traits, attention-based filtering was proposed in emotionally supportive dialogue systems. In this approach, structured user summaries were generated and used alongside dialogue content to modulate persona representations via attention mechanisms [62]. These architectures allowed AI agents to respond in emotionally appropriate ways while simultaneously inferring context-aware persona elements. Building on these foundations, in-dialogue learning (IDL) was proposed to enable the continuous inference of both static and dynamic personas during multi-turn interactions [63]. In IDL, dialogue history was segmented and analyzed to identify salient persona features, which were then incrementally updated throughout the session. Unlike prior systems that relied on fixed input or one-shot extraction, IDL allowed AI systems to revise their understanding of the user mid-dialogue without requiring predefined profile injection. To further enhance personalization, user engagement and subjective interest were modeled using neural signals during the interaction. In such models, latent cues such as emotion, surprise, or preference shift were tracked to inform real-time response adjustment [64]. By modeling engagement signals dynamically, these systems increased their sensitivity to subtle user behavior changes and improved their alignment with user expectations. Recent approaches have extended real-time preference adaptation using reinforcement learning and selection-based fine-tuning mechanisms. Adaptive preference learning strategies have been proposed to select the most suitable personalized output among multiple candidates, based on dialogue context and evolving user profiles [65], [66]. These models have allowed AI systems to refine their behavior mid-session, transitioning from reactive adaptation to proactive, user-aligned generation.
4.2.3 Post-Interaction Feedback and Preference Refinement
While predefined profiling and real-time preference detection enable short-term personalization, they remain limited in their ability to support long-term user alignment. In both prior phases, adaptation is constrained by the information available before or during the interaction. However, the outcomes of a completed conversation provide valuable signals that can be used to refine future behavior. As a result, preference detection techniques have evolved to include a post-conversation phase, where feedback from past interactions is utilized to optimize future responses.
Initial methods in this space were built around lightweight feedback integration, where human judgments were used to revise system outputs. To structure this process, human-in-the-loop frameworks were introduced to determine when intervention should occur and how user feedback could be mapped to corrective strategies. In one such approach, the appropriateness of system suggestions was judged based on relevance to prior user intent, and subsequent adjustments were made to improve alignment in future turns [66]. This method demonstrated how user feedback could guide post-hoc tuning of suggestion timing and content, particularly in co-creative or supportive dialogue environments. To generalize feedback across users, active preference learning systems were developed. In these frameworks, user preferences were elicited through minimal feedback signals, such as pairwise rankings or accept/reject decisions. These signals were then used to guide updates in the model’s preference space through reinforcement-like loops [65]. This strategy allowed systems to improve their personalization behavior over time, even when explicit supervision was unavailable. Compared to traditional tuning methods, these feedback-driven strategies required fewer labeled examples and facilitated continual learning from sparse interactions. A more principled foundation for feedback integration was formulated through models that incorporated preference alignment objectives at the optimization level. In one such method, preference optimization was performed using direct feedback criteria collected after interaction, enabling the model to adapt its internal persona representation based on past interaction quality [63]. This approach, referred to as DPOC (Direct Preference Optimization with Criterion), enhanced long-term alignment by assigning learning weights based on how well past dialogue aligned with persona-consistent objectives. To further extend alignment, frameworks such as AFSPP (Agent Framework for Shaping Preference and Personality) were proposed to enable continual preference shaping through long-term user engagement [67]. In these models, personality traits and user inclinations were refined across multiple sessions through a feedback loop involving both explicit evaluations and implicit user behavior. Unlike static persona models, AFSPP incorporated the gradual evolution of preference signals, leading to more adaptive and personalized AI agents. Ethical and conceptual implications of preference optimization were also explored, highlighting the tension between user satisfaction and long-term alignment goals. It was emphasized that feedback-driven systems must balance personalization with fairness and transparency, especially when user preferences evolve unpredictably or conflict with system goals [68]. These perspectives contributed to a more cautious and principled understanding of how feedback should inform preference modeling over time. Recent advances in retrieval-augmented systems have also demonstrated how generation quality can be improved through post-conversation evaluation. In one such architecture, response quality metrics such as BLEU and F1 were used to provide learning signals to the retriever, which was jointly trained with the generator [69]. This approach enabled the system to refine its input selection process based on the success of generated responses, leading to improved alignment in future interactions without requiring explicit human annotation.
Together, these models represent a significant evolution in preference detection, shifting from static or real-time adaptation to a reflective, feedback-driven process. By incorporating signals from completed dialogues, AI systems have been enabled to refine their personalization strategies over time, leading to more adaptive, consistent, and user-aligned behavior in future interactions.
Table 3: Summary of Preference Detection Techniques Across Conversation Phases.
Phase Techniques (Models/Methods) Personalization Focus Advantages Limitations
Before the Conversation Predefined profile modeling [54] Gaussian Processes [56] Session embedding [57] Static profiling Probabilistic modeling Dynamic session embeddings Initial preference modeling Profile construction Coarse personalization Population trends capturing Fixed user profiles Limited adaptivity Low contextual sensitivity
During the Conversation Zero-shot persona attribute extraction [61] In-dialogue learning [63] Neural Collaborative Filtering [60] Persona extraction Dynamic behavioral modeling Embedding-based preference adaptation Dynamic persona modeling Real-time context adaptation Fine-grained trait inference Latent behavior integration High model complexity Content-bound inference Computationally intensive
After the Conversation Feedback-driven persona optimization [63]; Continual personality shaping [67]; Active Preference Learning [65]; Feedback integration Long-term persona refinement Active learning from sparse feedback Long-term persona alignment Supports continual adaptation Efficient use of sparse feedback Requires labeled data Prone to feedback bias User privacy risks
4.3 Personalized Response Generation
Considering that the usable clues for preference optimization are now available through any of the previously discussed techniques, this section reviews different groups of approaches for response personalization. To this end, we have selected the techniques based on their technical diversity, rather than following a progression-based perspective on how this context has advanced over time.
Table 4 presents a range of response personalization studies applied at various levels of dialogue system generation. These include prompt-based approaches, which are model-agnostic, as well as more complex methods that propose entirely new architectures from scratch. In reviewing these studies, we considered a shared set of personalization aspects—such as consistency, coherence, and diversity—as identified in previous work [70, 71].
Personalization Level Core Mechanisms Personalization Focus Advantages Limitations
Prompt-Based Hybrid Graph-Augmented Prompting [72] Topic-Focused Summarization [73] RSA-Based Token Reweighting [74] Dual Retrieval Prompting [75] Selective Soft Prompting [76] Persona Consistency Persona Diversity Persona-Grounded Knowledge Dynamic Style Adaptation Model-agnostic Low computational cost Flexible to context Shallow persona integration Dependent on retrieval quality
Fine-Tuning-Based Emotion-Conditioned LoRA Fine-Tuning [77] Cue-Adaptive Decoding [78] Multi-task fine-tuning [79] Data-Centric Curriculum Fine-Tuning [80] Prefix Tuning [81] Interpretability Emotion Alignment Naturalness Fluency Memory-efficient Fast adaptation Low storage overhead Limited deep reasoning Risk of overfitting on small persona data
Architecture-level Concept Expansion and Filtering[82] Latent Act Modeling [83] Multi-Hop Split Memory Attention [84] Multi-Hop Co-Attention Matching [85] Addressee Prediction with Posterior Supervision [86] Persona-Adaptive Cross-Attention [87] Graph Transformer with Attribute Nodes [88] Deep Persona Grounding Multi-Party Dialogue Setup Dialogue Act Control Persona Coherency High expressiveness Fine-grained Reasoning High training cost Requires large labeled datasets
Table 4: Comparison of personalized response generation approaches organized by personalization level, mechanism, focus, advantages, and limitations.
4.3.1 Prompt-Based Personalization
As a straightforward class of methods, prompt-based approaches steer the model’s behavior by modifying the input through prompt engineering or retrieval-augmented generation (RAG), rather than altering the model itself. Their key advantages include model-agnostic applicability, low computational cost, and flexibility at inference time, making them well-suited for lightweight or rapidly deployable systems.
A variety of persona coverage and injection approaches have been proposed to carefully construct inputs and avoid altering the underlying language model. In CKE (Cross-Graph Knowledge Exchange) [72], a hybrid structured prompt is constructed by first generating dialogue user graphs for both conversation participants and then performing cross-graph knowledge aggregation. Discrete and continuous representations are fused to form a prompt that encodes persona, dialogue, and concept-level relationships for input to a frozen LLM. Similarly, a topic-focused summarization method [73] is applied to distill relevant context segments from the dialogue history. These summaries, combined with structured profile attributes and topic-aligned dialogue segments, are assembled into a prompt to guide generation in domain-specific tasks. Beyond structural augmentation, [74] employs token-level probabilistic reasoning using the Rational Speech Acts (RSA) framework. A self-conscious speaker adjusts token probabilities by simulating an imaginary listener, supported by a distractor memory storing persona variants, to enforce utterance distinctiveness and consistency during decoding. In PK-ICR [75], a dual retrieval prompting strategy is implemented by leveraging both dialogue context and persona to retrieve knowledge facts through a neural QA retriever, enabling precise grounding for generation. Lastly, Selective Prompt Tuning (SPT) [76] introduces a soft prompt selection mechanism, where multiple trainable soft prompts are maintained and a dense retriever selects the most suitable prompt per input. Contrastive and fusion learning objectives are further applied to ensure prompt diversity and stability across dialogue contexts.
4.3.2 Fine-Tuning-Based Personalization
While prompt-based methods are simple and model-agnostic, they offer limited control over model behavior. In contrast, fine-tuning-based personalization enables more robust and persistent behavioral adaptation, making it suitable when style adaptation or domain-specific alignment is required beyond what input conditioning alone can achieve.
To enhance persona consistency in generation, TransferTransfo [79] fine-tuned a pretrained Transformer using a multi-task objective that combined language modeling and next-utterance classification. Persona sentences were prepended to the input and processed through the model’s existing embedding layer, allowing the system to learn persona-sensitive representations without architectural modification. Cue Adaptive Decoder (CueAD) [78] introduced a decoder-side mechanism using GRU+, a modified GRU with a gating function that selectively incorporated conditioning cues, such as persona or dialogue context, based on their relevance during decoding. This enabled dynamic and interpretable control over cue influence. The D³ framework [80] took a data-centric approach, applying a three-stage fine-tuning pipeline: distillation to ensure alignment, diversification to expand behavioral coverage, and curriculum training to manage input complexity. This structured process helped improve generalization and consistency. In PersonaPKT [81], prefix tuning was adopted to personalize generation through continuous prefix vectors injected into a frozen backbone. A shared prefix was first trained across multi-persona data and then specialized into persona-specific prefixes, enabling private, modular adaptation. Emotionally adaptive generation was later explored by Varma et al. [77], where LoRA-based fine-tuning was applied to an LLaMA-3 model. An emotion classifier triggered prompt selection, guiding the generation of responses conditioned on emotional tone and demographic context. Extending this idea, Tan et al. [89] proposed a parameter-efficient method in which small, LoRA-based modules—referred to as pieces—are retrieved and composed at inference time. These modules enable storage-efficient, task-free personalization without altering the base model.
4.3.3 Architecture-Level Personalization
Despite the increased flexibility offered by fine-tuning methods, they remain bounded by the representational capacity of the underlying architecture. When personalization demands richer structural modeling, such as dynamically encoding speaker roles, incorporating graph-based knowledge, or explicitly handling multi-source inputs, modifications to the model’s architecture become necessary. At this level of intervention, personalization is achieved by introducing new components or redesigning internal mechanisms to explicitly encode and integrate user-specific signals.
Architectural personalization methods have progressively evolved to overcome the limitations of earlier prompt- and fine-tuning-based techniques, particularly in representing structured user information and modeling dynamic conversational behavior. One of the earliest efforts introduced discrete latent variables to represent conversational intent [83], allowing the model to generate responses conditioned on latent dialogue acts. This enabled one-to-many response modeling and helped align persona traits with underlying communicative goals. Building on the need to capture different dimensions of user information, memory-based architectures were proposed [84], where user profiles and interaction histories were encoded into split memory networks, and multi-hop attention mechanisms allowed for selective retrieval of static and dynamic user signals during generation. To better model interactions in multi-party settings, graph-based structures were introduced [86], encoding utterances, speakers, and dialogue flow into a heterogeneous graph, while also incorporating an addressee prediction module trained with posterior supervision. This line of work was extended by incorporating user persona attributes directly as graph nodes, connected via meta-relational edges for fine-grained reasoning [88]. Meanwhile, matching-based models such as CoBERT [85] leveraged multi-hop co-attention to capture second-order interactions between context, persona, and candidate responses, supporting persona-consistent selection in retrieval-based setups. Further developments introduced adaptive attention mechanisms within generative models [87], where cross-attention over persona and dialogue context was modulated through dynamic weighting and masking to filter irrelevant signals at each decoding step. In parallel, broader architectural pipelines were designed to address sparse or fragmented persona inputs, combining concept-based persona generation with alignment modules before feeding curated persona content into the generation process [82]. Together, these advancements reflect a shift from modular intent modeling toward structurally integrated architectures capable of deeply encoding and reasoning over rich user information within the generation pipeline.
4.4 Feedback-Driven Preference Optimization
In contrast to the approaches discussed in the previous section—which personalize responses by conditioning on explicit user preferences—another line of research focuses on post-training, preference-based alignment methods. These methods aim to align model behavior with human preferences by optimizing the model based on feedback over its generated outputs. A widely adopted approach in this category is Reinforcement Learning from Human Feedback (RLHF) [90], which involves training a reward model from human-labeled preferences and then fine-tuning the language model through reinforcement learning. In contrast, Direct Preference Optimization (DPO) [91] simplifies this process by eliminating the need for both a reward model and reinforcement learning. Instead, it directly optimizes a preference-based objective over response pairs using a supervised learning framework. In the following subsections, we review both RLHF and DPO in detail, highlighting their design choices, challenges, and recent advancements.
4.4.1 RLHF
While Supervised Fine-Tuning (SFT) can produce instruction-following behavior by training a model on labeled examples, it is inherently limited by the static nature of its training data, which does not reflect user preferences that arise during real-world deployment [90]. In contrast, Reinforcement Learning from Human Feedback (RLHF) augments this approach by introducing a post-training alignment phase, where the model learns to optimize for outputs judged more favorably by humans [92, 93]. Instead of relying solely on direct demonstrations, RLHF incorporates a reward model trained on human preference comparisons between multiple candidate responses. This reward signal guides further model updates via reinforcement learning [94], enabling it to adapt to subjective and nuanced human values, especially in open-ended tasks lacking a single correct answer [94, 95]. RLHF thereby extends beyond SFT by embedding a dynamic feedback loop that allows ongoing refinement of model behavior based on human-aligned criteria, improving its helpfulness, harmlessness, and honesty over time [96, 97].
Table 5: Summary of Key Focus Areas in the RLHF Pipeline. RLAIF stands for Reinforcement Learning from AI Feedback.
Category Primary Goal Key Techniques Main Challenges
Reward Modeling Preference estimation Feedback interpretation Pairwise comparison [98] Logistic heads [92] Reward ensembles [99] Dense feedback via credit assignment [100] Adaptive scaling [101] Overfitting Label noise Misaligned incentives
Policy Optimization Behavior alignment Actor-Critic methods [102, 103] KL-penalty regularization [104, 105] Rejection sampling [106, 107] Instability Reward hacking Exploration vs. alignment tradeoff
AI-Driven Feedback Feedback automation Human effort reduction RLAIF [108] Self-critiquing [109] Synthetic annotations [110] Bias propagation Hallucination risks Trustworthiness of AI-generated feedback
A central pillar of the RLHF pipeline is reward modeling, which serves as the mechanism for translating human preferences into a trainable signal that guides policy optimization. This approach was first formalized through pairwise preference learning, in which reward models were trained using binary comparisons between model outputs, as shown in Deep RL from Human Preferences (2017) [98]. Subsequent improvements extended this method to domains such as language generation [92] and summarization [93], where human preferences were systematically incorporated to better capture quality and alignment. As the field has matured, scalability and fidelity have become pressing concerns. To address computational limitations, ensemble-based reward modeling techniques were developed to enhance performance without excessive overhead [99]. To improve feedback quality, Attention-Based Credit Assignment was proposed as a mechanism for generating dense reward signals from sparse human input [100]. In parallel, the Adaptive Preference Scaling framework introduced a dynamic way to modulate human preferences in multi-objective settings, helping models align with more nuanced value signals [101]. Conceptual challenges—such as reward hacking and misgeneralization—have also been systematically outlined in foundational critiques of RLHF reward modeling [111]. Together, these efforts reflect a clear evolution from handcrafted reward structures to scalable, robust, and statistically principled models capable of capturing complex, subjective human values—firmly establishing reward modeling as a cornerstone of alignment in RLHF.
Policy optimization in RLHF is the stage where the learned reward signal is used to fine-tune the model’s behavior, often under alignment constraints. A common approach involves actor-critic frameworks, which enable sample-efficient updates by decoupling policy learning from value estimation. Recent work has extended this paradigm to constrained optimization settings to address reward overoptimization and instability in alignment objectives [102], and hierarchical actor-critic architectures have been proposed to manage multi-turn decision-making in complex language tasks [103]. To ensure the policy does not drift too far from its initial behavior, KL-penalty regularization is widely employed. While initially implemented as a global constraint, more recent advances have introduced token-level KL penalties that offer finer-grained control [104], alongside formal analyses that establish conditions under which KL regularization promotes both stability and alignment [105]. Another key technique, rejection sampling, has been explored as a safeguard against policy updates from low-quality feedback. Statistical rejection schemes have been proposed to filter unreliable preference comparisons, improving the fidelity of learning signals [106], while hybrid strategies like RS-DPO combine rejection filtering with direct preference optimization to enhance robustness in real-world deployment [107]. Collectively, these approaches illustrate how diverse optimization strategies are employed to balance safety, efficiency, and fidelity in RLHF-driven model alignment.
Recent research in RLHF has explored the use of AI-generated signals as alternatives or complements to direct human feedback, giving rise to the broader class of AI-driven feedback methods. In Reinforcement Learning from AI Feedback (RLAIF), preference data generated by language models is used to bootstrap reward models, reducing reliance on costly human annotation. This approach has shown promise in early experiments with self-generated comparisons guiding alignment [108]. A related line of work involves self-critiquing, where models learn to assess their own outputs or simulate preferences under uncertainty. Risk-aware methods have leveraged such self-feedback to refine policy training and improve robustness in alignment tasks [109]. Finally, techniques based on synthetic annotations aim to extract alignment signals from unlabeled or weakly labeled data sources. By leveraging proxy metrics or consistency within large corpora, these methods enable scalable preference modeling without explicit supervision [110]. Together, these techniques mark a shift toward more scalable and autonomous alignment strategies in the RLHF pipeline.
4.4.2 DPO
Although RLHF has proven effective in aligning large language models with human preferences, its reliance on reinforcement learning algorithms—often involving reward models, sampling loops, and KL-regularized objectives—can introduce complexity, instability, and inefficiency. To address these challenges, DPO has emerged as a simplified yet powerful alternative. Rather than learning a separate reward function and optimizing a policy through reinforcement learning, DPO directly fine-tunes the policy using pairwise human preference data. By framing alignment as a classification problem between preferred and dispreferred responses, DPO bypasses explicit reward modeling and policy rollouts, offering a more stable and scalable training pipeline. This paradigm shift not only streamlines implementation but also improves sample efficiency, making DPO an increasingly popular strategy for preference-based alignment in large-scale language models.
Variant Scoring Function $\mathcal{S}(y_{w},y_{l},x)$ Regularization Term $\mathcal{R}(y_{w},y_{l},x)$ Motivating Limitation Approach
DPO [91] $\beta\cdot\left[\log p(y_{w}\mid x)-\log p(y_{l}\mid x)\right]$ 00 Lack of explicit reward model Log-likelihood difference
CPO [112] $\beta\cdot\left[\log p(y_{w}\mid x)-\log p(y_{l}\mid x)\right]$ $-\log p(y_{w}\mid x)$ Overconfident updates Likelihood penalty
ORPO [113] $\lambda\cdot\left[\log\frac{p_{0}(y_{w}\mid x)}{1-p_{0}(y_{w}\mid x)}-\log\frac{p_{0}(y_{l}\mid x)}{1-p_{0}(y_{l}\mid x)}\right]$ $-\log p(y_{w}\mid x)$ KL-dependence Odds-ratio scoring
SimPO [114] $\frac{\beta}{|y_{w}|}\log p(y_{w}\mid x)-\frac{\beta}{|y_{l}|}\log p(y_{l}\mid x)-\gamma$ 00 Length bias Length-normalized scoring
IRPO [115] $\beta\cdot\left[\log\frac{p(c_{w},y_{w}\mid x)}{p_{t}(c_{w},y_{w}\mid x)}-\log\frac{p(c_{l},y_{l}\mid x)}{p_{t}(c_{l},y_{l}\mid x)}\right]$ $-\alpha\cdot\frac{\log p(c_{w},y_{w}\mid x)}{|c_{w}|+|y_{w}|}$ Poor generalization Compression + coverage regularization
$\beta$-DPO [116] $\beta\cdot\left[\log p(y_{w}\mid x)-\log p(y_{l}\mid x)\right]$, with adaptive $\beta$ 00 Static optimization sharpness Adaptive temperature
DPOC [63] $\beta\cdot\left[\log p(y_{w}\mid x)-\log p(y_{l}\mid x)\right]$ $P(r_{cho},r_{crt})+P(r_{crt},r_{rej})$ Preference misranking Criterion-based penalty terms
Table 6: Comparison of DPO variants by scoring functions, regularization strategies, motivating limitations, and optimization approaches. The DPOC regularization uses: $P(r_{a},r_{b})=-\min(0,\log r_{a}-\log r_{b})$. Table adapted from [2].
At the core of DPO lies a simple yet effective probabilistic framework that formalizes preference optimization without relying on explicit reward modeling. Given a pair of outputs—one preferred ($y_{w}$) and one rejected ($y_{l}$)—the objective encourages the model to assign higher likelihood to the preferred response relative to the rejected one. To support a unified view across emerging DPO variants, we generalize the original objective into the following form:
$\mathcal{L}(\theta)=-\log\sigma\left(\mathcal{S}(y_{w},y_{l},x)\right)+\mathcal{R}(y_{w},y_{l},x)$
Here, $\mathcal{S}$ is a scoring function that quantifies the preference between responses, and $\mathcal{R}$ is an optional regularization term. The original DPO formulation is recovered by setting $\mathcal{S}(y_{w},y_{l},x)=\beta\cdot\left[\log p(y_{w}\mid x)-\log p(y_{l}\mid x)\right]$ and $\mathcal{R}=0$. This closed-form structure captures the key intuition behind DPO: to directly align model behavior with human preferences in a stable, efficient, and extensible way.
Several algorithmic extensions to the original DPO formulation have been proposed to address practical limitations such as overconfidence, verbosity, prior dependency, and optimization sharpness. These are summarized in Table 6, which categorizes variants by their scoring and regularization formulations. To mitigate reward overoptimization and encourage more calibrated updates, Contrastive Preference Optimization (CPO) [112] adds a log-likelihood penalty term to the loss, discouraging the model from assigning excessively high probabilities to preferred responses. Odds-Ratio Preference Optimization (ORPO) [113], in contrast, replaces the standard log-likelihood difference with a log-odds scoring function derived from a reference model, thereby anchoring preferences relative to base policy probabilities and removing the need for explicit KL penalties. Simplex Preference Optimization (SimPO) [114] targets verbosity by scaling log-probabilities inversely with response length, effectively normalizing the preference score to favor concise outputs. Information-Ratio Preference Optimization (IRPO) [115] integrates auxiliary compression and coverage terms into both scoring and regularization components, enhancing robustness to preference noise and improving generalization under limited supervision. $\beta$-DPO [116] revisits the temperature parameter in the scoring function, replacing the fixed $\beta$ with either learnable or input-dependent temperature values to dynamically adjust alignment sensitivity across prompts. Most recently, Direct Preference Optimization with Criterion (DPOC) [63] addresses preference misranking by adding penalty terms based on criterion responses, ensuring that the preferred output meaningfully surpasses both the rejected and an intermediate-quality reference. To support controllable alignment across multiple preference dimensions—such as helpfulness, honesty, and harmlessness—multi-objective frameworks have been explored. One such approach is MODPO [117], which generalizes DPO to accommodate vectorized preference signals, enabling simultaneous optimization over multiple alignment axes without collapsing them into a single scalar reward.
5 CONCLUSION
This survey examined the role of preference optimization in the design of AI copilots—AI-powered, context-aware systems that support users in real-time, high-context tasks. We began by establishing a unified definition of AI copilots that consolidates diverse interpretations across domains, providing a conceptual basis for consistent study and development. Building on this, we offered a phase-based view of preference optimization, structured around how user preferences are identified, interpreted, and used to drive personalized interaction throughout the lifecycle of user engagement. By synthesizing methods for detecting preferences, generating contextually appropriate responses, and refining behavior through feedback, this work provides a coherent framework for understanding and advancing personalization in AI copilots. Beyond conceptual synthesis, the survey offers practical guidance for system designers by mapping available preference sources, outlining how they can be integrated, and analyzing technique selection across each stage of a preference-aware AI copilot.
As a future direction, this perspective could be extended by systematically exploring implementation pathways and cost trade-offs. Each stage involves choices—such as platform selection, model customization, or data infrastructure—that carry implications for scalability, latency, and operational cost. Studying these dimensions would help with more actionable strategies for building effective and efficient preference-aware AI copilots.
Acknowledgments
We acknowledge the Centre for Applied Artificial Intelligence at Macquarie University (Sydney, Australia) for funding this research.